If g(x) = 2x - 3, D = {Reals}, and h(x) = x squared + 1, D = {Reals}, find:
a) (g + h)(-2)
b) gh(1/2)

Answer by mukhopadhyay(490)

(g + h)(x)= 2x-3+x^2+1 = x^2+2x-2
(g-h)(-2) = 4-4-2 = -2;
gh(x)= (2x-3)(x^2+1);
gh(1/2) = (1-3)(1/4 + 1) = -2(5/4) = -5/2